### v2

# spacing=nonfrench

Abstract
Specializing systems to specifics of the workload they serve and platform they are running on often significantly improves performance. However, specializing systems is difficult in practice because of compounding challenges: i) complexity for the developers to determine and implement optimal specialization; ii) inherent loss of generality of the resulting implementation, and iii) difficulty in identifying and implementing a single optimal specialized configuration for the messy reality of modern systems.
To address this, we introduce Iridescent, a framework for automated online system specialization guided by observed overall system performance. Iridescent lets developers specify a space of possible specialization choices, and then at runtime generates and runs different specialization choices through JIT compilation as the system runs. By using overall system performance metrics to guide this search, developers can use Iridescent to find optimal system specializations for the hardware and workload conditions at a given time. We demonstrate feasibility, effectivity, and ease of use.
1 Introduction
Specializing system implementations to workload characteristics and hardware can significantly improve performance and efficiency [27, 39, 82, 53, 78, 4, 41, 70, 31, 86, 30, 24, 72, 8, 55, 60, 77, 22, 18, 16, 76]. Achieving these benefits requires manual modification of the system implementations and recompilation. Part of the performance benefit arises from cascading compiler optimizations, e.g. by removing a feature enabling the compiler to eliminate dead code, in turn enabling further optimizations  [83, 61, 27]. Today, system specialization for performance is manual, developer-driven, and iterative.
Building optimal systems in practice is a challenge because of three compounding factors: First, specialization is difficult for developers to implement, as it requires trial-and-error and maintaining multiple versions of key code paths. Next, specialization fundamentally comes at the cost of generality — a specialized system either performs poorly outside its regime, or completely fails. Finally, predicting performance implications of specialization choices for different workloads and hardware is almost impossible. To make matters worse, workload and platform conditions are dynamic and optimal specialization choices depend on them.
We propose a fundamentally different approach: automated online system specialization guided by observed overall system performance. Our goal is to specialize performance critical system systems at runtime to the exact momentary workload and hardware conditions, without the human developer on the critical path. To enable this, we re-distribute tasks between ahead of time development and runtime operation. Ahead of deployment, rather than choosing a concrete set of specializations, developers specify the space of possible specializations, along with a search strategy. After deployment, our runtime iteratively generates and measures the performance of different specialized code versions, choosing the currently optimal specialization based on observed overall system performance. This allows compile-time specializations at runtime, suitable for search-based auto-tuning [82, 78, 75].
To enable this, we introduce Iridescent, a framework for practical development of fast and efficient systems with online specialization. Iridescent enables both the incremental specialization of existing code-bases with minimal modifications, and in-depth design for specialization for maximizing performance. Iridescent builds on LLVM and targets systems written in typical (non-managed) languages such as C(++) or Rust that are architected around an event-loop. Concretely, Iridescent requires developers to separate the code-base into a specialized, performance critical handler core, and the remaining outer event loop code. Iridescent then provides the developer with a specialization API to annotate possible specializations in the code thereby specifying the space of possible specializations. Additionally, Iridescent provides hooks to optionally customize JIT code generation in depth. Finally, Iridescent provides a runtime API to control the exploration of different specialized implementations as the system runs.
[…]
2 Motivation & Challenges
We illustrate the benefits and challenges of specialization with an example system: a server executing square matrix multiplications for clients. To optimize cache locality we use a blocked matrix multiplication algorithm [11, 84, 85]. We identify one key workload parameter, $N$, the matrix size, and a key configuration parameter, $B$ the block size.
Code specialization improves performance.
Both parameters offer a substantial opportunity for optimization through specialization. Unsurprisingly, we find fixing $B$ as a compile-time constant instead of leaving it a variable improves throughput by up to $6.5\times$, by enabling the compiler to unroll and vectorize the inner loops. Note that $B$ is an internal parameter that may affect performance, but any valid choice results in correct behavior with every workload. Similarly, assumptions about the workload, $N$, can also simplify the algorithm, e.g. by assuming that $N$ is a multiple of $B$ we avoid the need for copying and 0-padding partial blocks.
Machine / Workload N=1024 N=256 N=64
IceLake 32 32 32
IvyBridge 16 16 4
CoffeeLake 32 4 4
AlderLake-p 32 4 2
AlderLake-e 64 4 4
Table 1: Optimal value of block size, $B$, for our block matrix multiply, across 5 hardware platforms and 3 workloads.
Optimal configuration depends on workload and HW.
We now compare different block sizes for different workloads (matrix sizes) on 5 different processors. As Table 1 shows, different block sizes yield optimal performance for different workload and processor combinations. Picking a fixed $B$ ahead of a practical deployment in a dynamic environment will not yield optimal performance. Moreover, even with single single-size workloads it is difficult to predict what block size will be ideal for a concrete processor.
2.1 Specialization is Effective but Complex
There is a long line of prior work that has established the performance and efficiency benefits of specializing code, and systems code more specifically. For example, interpreted languages may generate specialized instances of functions with constant parameters and optimize them accordingly [21, 49]. This is a generic optimization and typically done automatically and transparently by the runtime. Other optimizations are specific to individual systems and concrete concerns, such as inlining small table lookups in software network functions as if statements [53]. Like most compiler optimizations, these approaches are typically guided by simple heuristics around local metrics (e.g. frequency of the same parameter value, or table size) based on developer intuition.
Specialization effects often cascade.
A key aspect of specialization techniques is that the effects combine. For example, PacketMill [27] first de-virtualizes function pointers [43] for the Click modular router [44], and then based on this further eliminates dead code and data structure fields. The analysis for the latter optimization is impossible before de-virtualization as the virtual calls prevent the compiler from analyzing the complete packet handling code.
2.2 Specialization is Challenging in Practice
(C1) Developer Complexity.
A key barrier to specialization is the increased complexity for developers. Developers now need to reason about what assumptions will definitely hold and can help specialize the system for better performance. Next the developer needs to implement the required specialization and evaluate it. In practice, this frequently results in having to maintain multiple code versions. Unsurprisingly, this is not only complex but also laborious and frustrating, since it is often an iterative hit-and-miss process.
(C2) Loss of Generality.
Specializations may hurt performance or break the system when the underlying assumptions are violated. For example, a network function may process 99.999% of packets if the processing loop and data structures are specialized in a specific way. However, specializing the system to this class of packets, and thereby simplifying code and data structures may break handling of the rare 0.001% cases. The need to gracefully handle rare or unexpected cases forces developers to specialize conservatively.
(C3) Optimal Specialization Choices.
Modern system performance is a complex product of the emergent behaviors across all system components, it is also not feasible for developers to distill this into simple local cost models. Optimally specializing systems statically is fundamentally impossible since concrete workload and hardware parameters affect these choices, and both are dynamic in practice.
3 Approach
Specialization Effect
Value Convert runtime variables to constants; triggers const. prop., unrolling, vectorization, ….
Assumption Add boolean assumptions; triggers dead code elimination, branch elimination, …
Fast-Path Add if-else based early exits with previously computed/cached results.
Custom Code Generate custom code at runtime. e.g. dynamic LPM matching via nested branches.
Table 2: Iridescent specialization types and their effects.
We enable practical specialization of systems for improved performance and efficiency. Our key idea to enable this given the challenges above, is to automatically explore a space of possible specializations provided by the developer at runtime based on observed overall system performance.
Simple specialization with simple annotations.
We first observe that, while developers struggle with determining optimal specializations and implementing a system based on them, it is much easier for developers to suggest possible specialization assumptions. We thus enable developers to annotate system code with simple annotations to this end. 2(a) shows an example of the annotations for the matrix multiplication server above and how developers can configure specializations at runtime. Additionally, developers can also provide LLVM transformations to perform intrusive app-specific specializations, and thereby generate further candidate configurations in the overall specialization space. Crucially, we do not rely on the developer to filter or rank specializations. Table 2 shows the broad class of specializations supported by Iridescent and their corresponding effects.
Guarded specialization with a fall-back.
Later, when automatically instantiating specialized system versions through JIT compilation, we inject guard conditions into the code to use the specialized version of the code when applicable. If the specialized version is not applicable, the guard conditions ensure that we fall back to the unspecialized version of the system code for that invocation. While triggering this guard incurs overhead, it does ensure correctness. If the guard triggers sufficiently rarely, in combination with the benefit for the common case, the specialization is still a net win.
Online exploration guided by developer policy.
Finally, we completely forgo cost models and heuristics for predicting which choice in the specialization space is optimal, and instead explore different points online as the system runs. Our guards ensure that the system always behaves correctly in this process. We rely on the developer to provide a policy for guiding this exploration. The developers control when to explore which points, manually, through existing auto-tuning solutions or by leveraging simple search strategies from our library. A key task of this policy is comparing the (application specific) overall system performance metrics, e.g. request throughput, when running different specialization points. Overall system performance metrics also implicitly factor in any overheads, e.g. triggering specialization guards for some calls. This ensures that the system converges to the optimal specialization point for the concrete combination of dynamic system, workload, and hardware conditions.
4 Iridescent Design
We now present the design of our approach in Iridescent.
Figure 1: Architecture of a Iridescent-specialized system.
1void matmul(u64 *L, u64 *R, u64 *O, int N, int B) {
2 B = spec_enum("B",B,2,4,8,16,32,64);
3 N = spec_general("N",N);
4 for (int $k$ = 0; $k$ < N; $k$ += B)
5 for (int $j$ = 0; $j$ < N; $j$ += B)
6 for (int i = 0; i < N; i++)
7 for (int jjj = $j$; jjj < $j$ + B; jjj++)
8 for (int kkk = $k$; kkk < $k$ + B; kkk++)
9 M3[i*N+jjj] += L[i*N+kkk] * R[kkk*N+jjj];
10}
(a) Handler Code
1Iridescent rt("handler_code.ll");
2void spec_policy() {
3 Conf N_cs[2] ={{}, {{"N", 256}}};
4 do { for (c : cartesian(rt->spec_space(), N_cs)) {
5 rt.specialize(c); tput_counter = 0;
6 sleep($\ldots$); // Sleep for some time to monitor
7 if (tput_counter > best) {
8 best = tput_counter; best_c = c; } }
9 rt.specialize(best_c);
10 } while(every 1min);
11}
12int main() {
13 auto *matmul = rt.handler("matmul");
14 rt.reg_opt_pipeline(opt_passes);
15 launch_thread(spec_policy);
16 while(true) { // Main processing loop
17 Req *req = get_req(); Resp *res = new Resp();
18 matmul(req.A, req.B, res.C, 2, req.N);
19 tput_counter++;
20 }
21}
(b) Fixed Code
Figure 2: MMulBlockBench microbenchmark
Iridescent Specialization API
[…]
spec_generic(lbl, x) Policy-controlled spec. point
spec_assume(lbl, cond) Specialization assumption
spec_custom_*(lbl, ...) Custom spec. point
Iridescent Policy API
Iridescent(handler_ir) Initialize runtime
.handler(h) Get specialized handler
.spec_space() Obtain code spec. space
.specialize(c) Choose specialization; c maps spec. point labels $\mapsto$ values
.add_custom_spec(n, gen) Add app specialization
.customize_opts(passes) Modify codegen optimizations
Table 3: Iridescent API overview.
4.1 Anatomy of a Iridescent System
System Requirements for Iridescent.
Iridescent targets performance critical systems not implemented in managed languages. We observe that typical performance critical systems code is almost exclusively architected with an event loop structure – an outer loop handling a sequence of events, such as processing requests or data elements, and an inner handler that processes the event. While Iridescent is designed to specialize these systems in varying dynamic environments, for these systems to fully gain the benefits of Iridescent, these systems should have stability in the workload and operating conditions for a sufficient period of time. Conditions should at least remain stable for the time Iridescent takes to re-compile a new specialized variant (typically 10–400 ms, see § 6.4).
Isolating performance critical code for specialization.
Since Iridescent targets performance critical systems not implemented in managed languages, modifying code of the running system is non-trivial in general. We leverage the event-loop structure for pragmatic specialization, by requiring developers to separate the system code into two parts (Figure 1): (i) the fixed code, including initialization and the outer loop; (ii) the performance-critical event handler code. The developer compiles the fixed code as before, but integrates calls to the Iridescent API (Table 3). In the handler code, the developer adds lightweight Iridescent specialization API calls and compiles to LLVM intermediate representation. In the fixed code, the developer obtains function pointers for the handlers through the Iridescent policy API, and calls these as before. Figure 2 illustrates this with our matrix multiplication server running example.
Handling Global State.
During dynamic code generation, Iridescent will link handlers against symbols in the fixed code, to enable calls and accesses to global state. Since Iridescent repeatedly rebuilds handlers, Iridescent requires that the handler code does not include definitions of global variables that must be preserved. The developer has to move all global state to the fixed code, and reference it from the handlers.
4.2 Defining Specialization Space
The developer defines the specialization space through specialization point annotations in the handler code. At runtime, Iridescent replaces these annotations with specialized code, instrumentation, or disables them, guided by the policy. Iridescent offers four types of specialization points:
Value Specialization Point.
A value specialization point signals to Iridescent that the handler code could be specialized to a constant for the wrapped expression. spec_enum indicates the value could be one of the specified values; spec_range instead specifies a range of integers; spec_generic leaves it to the policy to determine possible values.
Assumption Specialization Point.
This provides a possible specialization assumption to Iridescent. The condition is a boolean predicate that Iridescent can provide to compiler optimizations to further optimize, through LLVM’s builtin llvm.assume. Crucially, the assumption need not always hold. Unlike for the LLVM intrinsic where incorrect behavior results, Iridescent catches this with its specialization guards.
Fast-Path Specialization Point.
This is a generic version of the hot key specialization of Morpheus [53]. It works in two phases: (i) the instrumentation phase: Iridescent inserts instrumentation code that samples (sampling rate selected by the developer) invocations of the specialized function to find the most popular inputs to the function along with their calculated outputs; (ii) the specialization phase: Iridescent updates the specialized function to have a specialized if-else chain where the top-N inputs along with their outputs are converted into a series of if-else checks to avoid paying the cost of re-computation for the heavy-hitting inputs. The value of N may be developer-provided or could be configured as a runtime constant that can be further specialized. If the input does not match any of the branches in the if-else chain, the computation simply defaults to the generic version.
Custom User-Defined Specialization Point.
These specialization points, annotated through spec_custom_*, invoke developer-defined specializations registered in the fixed code through add_custom_spec. An example is generating an if-else chain as a special-case fast-path for an expensive data structure lookup or computation (§ 5.2).
4.3 Defining Specialization Policies
The developer implements the specialization policy in the system’s fixed code using the Iridescent policy API. The policy decides what, when, and how Iridescent should specialize.
2(b) illustrates how developers can control the exploration and selection of specializations. In line 4, the spec_space API returns the specialization space generated by the annotations in the handler code. We combine this space with other specializations through a Cartesian product to obtain a complete set of specializations. In lines 4–8, the fixed code automatically tries out the different combinations in the set of specializations and chooses the best performing combination. To choose the best, the specialization policy uses the target end-to-end performance metric, in this case the throughput. We trigger re-exploration at a fixed-time interval to adapt to workload changes.
The developer can also optionally configure custom optimization passes through the customize_opts API to interpose on and customize code generation. Developers may also choose to enable instrumentation for select specialization points to dynamically identify opportunities for specialization. For example, the policy for the matrix multiplication server could detect the most frequent matrix size $N$, and only decide to specialize if $\geq$70% of requests have this value.
Exploring the specialization space.
Iridescent provides a periodic exhaustive search strategy for simple cases as a library routine, with a simple call-back for providing the strategy with the system performance metric. For more complex cases, Iridescent integrates with OPPerTune [75] to leverage its state-of-the-art search strategy for configuration tuning of web services. Additionally, Iridescent also provides a thin interface for developing custom search strategies. We expect most systems will either utilize existing auto-tuning or implement custom strategies maximizing overall system performance, using the Iridescent building blocks.
4.4 Specialization Runtime
The specialization runtime has two key components: (i) the specializer generating code by applying selected specializations; (ii) the JIT compiling and optimizing the modified specialized code and making it available to the fixed code.
4.4.1 Specializer
The specializer generates the specialized version of the handler code. It operates at function granularity, and generates versions of functions with specialization points replaced depending on the chosen specialization configuration.
For each specialization point, Iridescent performs the following actions, depending on the configuration the policy supplied for this point. If the policy marks a point as disabled, the Iridescent specializer simply removes the annotation and replaces it with the original value, or skips over the point if it is an assumption point. To specialize an enabled value specialization point, Iridescent replaces the specialization point annotation in the handler code with the constant value supplied by the policy. By default, the specializer will also insert a specialization guard, which the developers may explicitly disable. For a custom specialization point, the specializer replaces the specialization point with the custom source code provided by the previously registered handler for that point type. The recompilation process then enables standard compiler optimization to take advantage of the newly inserted constants or code.
Instrumentation.
Some specialization policies benefit from collecting runtime data to generate possible specialization configurations. To support such data collection, the specializer can optionally enable instrumentation for each specialization point. For a specialization point with instrumentation enabled, Iridescent additionally generates code for collecting and storing the observed actual values. The policy retrieves this information included in the result of the spec_space call.
4.4.2 JIT
Compiling the specialized code.
The specialization runtime adds the specialized code generated by the specializer to the JIT. The JIT compiles the code and runs the developer-specified pipeline of optimization passes, including any custom developer-provided optimization passes, to generate an optimized version of the specialized code. By default, the JIT applies the default O3 optimization pipeline if the developer has not provided a specific pipeline. In addition, the JIT also keeps a copy of the original, non-specialized version of the specialized code. This generic code serves as fallback for situations where the specialized code is not applicable.
Using the specialized code
. To use the specialized handlers, the fixed code obtains function pointers to specialized code from the specialization runtime. The fixed code then uses these pointers to invoke the specialized handlers. If no specialization has been enabled the handlers default to the non-specialized versions of the functions. For simplicity, the fixed code only does this once at start of execution. The JIT creates a trampoline function which calls the most recent specialized version of the function. The trampoline function is stored at a fixed address and does not change across specialization iterations.
4.4.3 Correctness through Specialization Guards
Iridescent inserts a runtime check called a specialization guard for each enabled specialization point. The policy can disable this guard for individual points. The specialization guard ensures that the specialization conditions for generated code holds during execution. For this, the specializer inserts an early exit from the specialized version of the code on guard failure by throwing an exception. The Iridescent trampoline function catches these exceptions and re-routes control to the original non-specialized version of the function transparently without exposing the exception to the fixed code.
Restoring state and side effects.
To ensure correct restoration of the state, the specializer will call a developer-defined cleanup function for reversing any side-effects before throwing the exception. It is critical to note here that not all side-effects are reversible (e.g. sending a packet to a neighbor), so Iridescent leaves the developer with the necessary control.
4.5 Prototype Implementation
We have implemented Iridescent in $5.5K$ lines of C++. Our implementation uses the LLVM IR and JIT [45] to generate the specialized code at runtime. The specializer is implemented as a set of LLVM transformation passes that operate on the LLVM IR of the instrumented handler code.
5 Case Studies
We evaluate Iridescent with four open-source systems and libraries and the MMulBlockBench microbenchmark. We explain each system along with the changes we made for integrating Iridescent below.
5.1 Tiling Benchmarks
MMulBlockBench.
We implement a blocked/tiled version of matrix multiplication algorithm [11, 84, 85]. To integrate this benchmark with Iridescent, we convert the tile size, $B$, as a value specialization point.
CloverLeaf.
We use a C-version of the CloverLeaf application from the SPEC ACCEL2023 benchmark [19]. In this application, we implement tiled-versions of the various mathematical kernels, and use value specialization points for the tile sizes, $T_{i}$ and $t_{J}$. In total, this application has a 24 unique specialization points, each of which that can take 7 unique valeus resulting in a total of $7^{24}$ possible configurations.
5.2 LibLPM
LibLPM [66] is an open-source Longest Prefix Match (LPM) library written in C with built-in support for IPv4 and IPv6 addresses. LPM is a key operation in packet routing for finding the closest matching route in routing tables for incoming packets. To integrate LibLPM with Iridescent, we create two different specializations:
LibLPM-FP
specializes the lookup function to add a fast-path specialization point. This specialization adds an exact match if-else fast-path for the top-N input IP addresses.
LibLPM-NI
creates a code-generation specialization point in the LPM library. The code-generation specialization point generates a new version of the lookup function which generates a nested-if-else tree of checks consisting of prefix match checks for the incoming address for each lpm entry. The nested-if-else tree starts at the least specific matching rule (based on prefix length) and progressively checks for the most specific matching rule until it can’t find one anymore. With this code generation specialization, we embed the prefix rules directly into the codebase, forgoing expensive dynamic datastructure traversals and enabling further optimizations for individual branches.
LibLPM-NI-FP
combines the previous two specializations by adding a fast-path specialization point in the generated nested-if specialized lookup function.
5.3 TAS
TAS [42], TCP acceleration as a service, is a lightweight software TCP network fast-path that is optimized for common-case client-server RPCs. TAS executes common-case TCP operations in an isolated fast path that uses DPDK [25], while handling corner cases in a slow path.
To integrate TAS with Iridescent, we convert its BATCH_SIZE macro to a value specialization point. In TAS, the BATCH_SIZE variable is used in three different scenarios: (i) to determine the number of packets to read from the NIC, (ii) to determine the number of packets to read from the application queues, and (iii) to determine the number of packets to read from the queue manager. For more fine-grained control, we convert each of these usage instances of the BATCH_SIZE to three separate runtime constant specialization points - (i) rx_batch, (ii) queues_batch, and (iii) qman_batch.
5.4 FastClick
FastClick [7] is an extended version of the Click Modular Router [44] with Netmap and DPDK support for running the modular router in userspace. A FastClick (or Click) router is assembled from individual packet processing modules called elements. Each element implements simple router functions such as packet classification, scheduling, routing, and interacting with network devices. Users configure the router with different graphs of connected elements.
To integrate FastClick with Iridescent, we modify the LinearIPLookup element of FastClick to create a Iridescent-enabled LinearIPLookup element. The original element uses a linear search algorithm to find the best matching route for an incoming packet based on the routing table. We modify the packet-processing function of the LinearIPLookup element to add a fast-path specialization point. The target of this specialization point is the internal lookup function of the LinearIPLookup element.
5.5 Network Functions
Network Functions is a suite of the network functions that include DPDK and ebpf implementations of common network functions such as a NAT, Router, policer, among others. These functions have been developed and used for evaluation by various research projects such as Pix [36] and Vigor [88]. We extract the implementations from the Pix artifact [35].
To integrate the Network Functions with Iridescent, we convert the BATCH_SIZE used by the DPDK version of these functions into a runtime constant specialization point. In the network functions, the BATCH_SIZE controls the number of packets that are read from a network port at a given time.
6 Evaluation
In this section, we evaluate how much Iridescent improves performance of systems by enabling online specializations. We show how we can use online specialization with Iridescent for three use-cases: (i) enabling compile-time optimizations at runtime; (ii) enabling incremental specializations at runtime; (iii) design exploration at runtime. We answer the following questions:
  • Do Iridescent-enabled optimizations improve performance? (§ 6.2)
  • Can Iridescent find an optimal specialization through exploration at runtime under dynamic conditions? (§ 6.3)
  • What is the overhead of using Iridescent? (§ 6.4)
6.1 Experimental Setup
Testbed.
For our experiments, unless otherwise stated, we configure two identical machines as client and server. They are directly connected with a pair of 100 Gbps Mellanox ConnectX-5 Ethernet adapters. Each machine has two Intel Xeon Gold 6152 processors at 2.1 GHz, each with 22 cores for a total of 44 cores and 187 GB of RAM per machine. We run Linux kernel 5.15 with Debian 11.
6.2 Compiler Optimizations at Runtime
Machine Constant (c) Variable (v) Benefit
(Processor) (cycles/op) (cycles/op) (v/c)
IceLake 175297 284944 61%
IvyBridge 250434 661295 246%
CoffeeLake 168817 581130 348%
AlderLake-p 173350 583724 336%
AlderLake-e 206924 557572 269%
Table 4: Impact of turning block size, $B$, as compile-time constant at runtime for N=64 for different hardwares (Table 1)
Figure 3: Average cycles required to perform a table lookup with varying table sizes under two different workloads.
Compile-time optimizations often produce more performant code. These compile time optimizations include constant propagation, loop unrolling, dead-code elimination, and use of vectorization instructions. With Iridescent, we can enable compile time optimizations using runtime data.
Iridescent enables cascading compiler optimizations.
First, we show the potential benefits of enabling compiler optimizations at runtime for the MMulBlockBench microbenchmark in Table 1. Specifically, for the workload $N=64$, we compare the performance of keeping the optimal block size, $B$, as a runtime parameter to that of converting it into a constant at runtime using Iridescent, thus enabling downstream compiler optimizations. For both configurations, we execute the function a fixed number of times and measure the number of cycles spent per function execution. Table 4 shows the benefit of converting $B$ into a compile-time constant for different machines. For each machines, we get at least 50% reduction in cycles, and greater than 240% reduction in cycles for four of the five operating conditions. This improvement is a direct impact of Iridescent’s specialization allowing different compiler optimizations to cascade and combine together to produce a more efficient version of the code. In this specific case, the specialization of $B$ allows the JIT compiler to first easily unroll the loops in the matrix multiplication function and then leverage optimized vector instructions.
Figure 4: FastPath router throughput with Iridescent specialized LinearIPLookup element for different routing table sizes at 100% fast-path hit rate.
Figure 5: FastPath router throughput with Iridescent specialized LinearIPLookup element for varying fast-path hit rates.
Iridescent incorporates existing specializations.
[…]
Iridescent enables custom specializations.
Next, we show the benefits of custom compile-time optimizations for LibLPM. For this experiment, we set up eleven different configurations resulting from a combination of five different LPM table sizes and two different workloads - Workload A and Workload B. In Workload A, the incoming IP address is an IP address that matches a very specific prefix entry in the routing table. In Workload B, the incoming IP address is an IP address that only matches the LPM prefix entry of prefix length 0. Thus, Workload A and Workload B cover the best and worst cases respectively for the LPM lookup function. We execute these workloads with seven specializations:
  1. (i)
    Baseline: no Iridescent-enabled specialization.
  2. (ii)
    Instrumented: Iridescent-enabled specialization capturing the most popular inputs (10% sampling); data used to guide the fast path specializations.
  3. (iii)
    LibLPM-FP-A: fast-path specialization of size 1 specialized for Workload A.
  4. (iv)
    LibLPM-FP-B: fast-path specialization of size 1 specialized for Workload B.
  5. (v)
    LibLPM-NI: nested-if code generation specialization.
  6. (vi)
    LibLPM-NI-FP-A: Combine LibLPM-NI and -FP-A.
  7. (vii)
    LibLPM-NI-FP-B: Combine LibLPM-NI and -FP-B.
Figure 3 shows average cycles per LPM lookup lookup function for all different combinations of configurations and specializations. For Workload A, all three specializations, LibLPM-FP-A, LibLPM-NI, LibLPM-NI-FP-A, achieve up to $6\times$ reduction in cycles per execution. However, unlike the LibLPM-FP and LibLPM-NI-FP-A specialization, this LibLPM-NI specialization does not require an additional instrumentation phase to sample executions to find the optimal candidates for the fast path. LibLPM-NI improves performance across the board for all workloads by up to $6\times$. Despite this performance boost, LibLPM-NI is not the best specialization for Workload B. For Workload B, the best optimizations are the *-FP-B specializations as they improve by up to $30\times$ as all instantiations of Workload B hit the fast-path.
6.3 Runtime Design Exploration
Figure 6: Automatic exploration and specialization of different configurations for TAS.
Iridescent enables runtime design exploration.
We show runtime design exploration for TAS [42]. We experiment in a typical server-client setting, with the server providing an echo service. The client is multi-threaded and executes an open-loop workload of 64 byte packets. Both the server and client are TAS-enabled. In our experiment, we modify the server-side TAS and measure the throughput of the server side TAS as millions of packets processed per second. With Iridescent we explore different values of the rx_batch in the TAS source code on the server side. Figure 6 shows Iridescent automatically determining and then selecting the best-configuration on the server side by exploring the different values of the batch size for different code locations.
Figure 7: Automatic exploration and specialization of MMulBlockBench for two different workloads.
Iridescent automatically adapts to changing workloads.
To show that Iridescent can automatically adapt to workloads, we first use the MMulBlockBench microbenchmark. We execute a Iridescent-enabled version of the function and a non-Iridescent enabled baseline. In the baseline, the block size $B$ is a runtime parameter. In the Iridescent-enabled version, Iridescent explores different values of $B$ by specializing the value in the code and converting $B$ into a compile-time constant. We execute each version with two different workloads in succession, each lasting one minute. We measure the overall throughput (executions/s) for both configurations. Figure 7 shows how Iridescent explores the different block sizes and finds the most performant configuration compared to the baseline. Moreover, Iridescent can automatically detect the workload change based on the drop in throughput and restart the A/B testing process.
Figure 8: Batch size exploration with Network Functions
Next, we show runtime design exploration and adaptation for two different network functions, vignat and vigpol, from vigor [88]. We setup our experiment with a packet generator on one machine and a device under test (DUT) on the other machine. We use two ports for the DUT. The first port acts as the internal facing port whereas the second port as the external facing port. The DUT executes the specific network function for every incoming packet generated by the packet generator. The network function implementation has different execution pathways for packets depending on whether they arrive on the internal port or the external port. The network function is enabled with the Iridescent specialization policy to trigger an exploration phase to find the best performing BATCH_SIZE whenever there is a change in the workload. We execute the experiment in two phases: in the first phase, packets arrive only on the external facing port and in the second phase, packets arrive only on the internal facing port. Figure 8 shows the results for the NAT and the Policer network functions. Iridescent finds the best performing configuration in each phase with different optimal values.
Figure 9: Optimal fast path size exploration with Iridescent.
Iridescent transparently enables exploration-based adaptation with custom specializations.
We show runtime design exploration for FastClick LinearIPLookup. We run the FastPath router on one machine, treat that machine as the Device Under Test and run Pktgen to execute an open-loop workload generating packets with destination ip addresses from a given set of IP addresses. At the 1 minute-mark, we completely switch the destination IP address set with no overlap with the initial address. The router is configured with a Iridescent specialization policy that triggers an exploration whenever it detects a large change ($\geq 25\%$) in measured throughput. Figure 9 shows how the throughput of the router changes for the different IP sets. At the start, the specialization policy triggers an exploration as it detects a large in-flux of packets, a significant change from 0 incoming packets. First, Iridescent runs a quick instrumentation phase for around 100ms to find the most popular incoming destination IP addresses. After the instrumentation phase, Iridescent switches to an exploration phase to explore the size of the fast-path, i.e., how many addresses should be included in the generated if-else source code. Once the exploration finishes, Iridescent switches to the best performing fast-path size and generates the specialized code which persists until the workload switch at the 1-minute mark. Due to the workload change, the specialization policy kicks in automatically, and Iridescent restarts the instrumentation and exploration phase. As shown in Figure 9, Iridescent can quickly adapt to changes in workload to generate optimal specialized code under dynamic conditions.
Iridescent can explore large specialization spaces.
We explore the large specialization space of CloverLeaf with OPPerTune’s exploration algorithm integrated with Iridescent. We limit the number of configurations to be explored to 20 out of a total of $7^{24}$ total configurations. We observe that, Iridescent can find an configuration that reduces the number of cycles spent per iteration of the main CloverLeaf loop by $3\times$ on average. By leveraging state-of-the-art auto-tuning exploration strategies, Iridescent can efficiently find an optimal configuration from large specialization spaces.
6.4 Iridescent but at what Cost?
Compilation cost.
Table 5 shows the JIT compilation time for each of the target systems. Note that, this compilation happens off the critical path so it is not a performance bottleneck. However, this cost does dictate the propagation delay, i.e., the time taken for the specialized version of the code to be available for the execution.
System Time (ms) Size (KB)
MMulBlockBench 10 $\pm$ 1 10
LibLPM-FP 72 $\pm$ 9 80
Network Functions 98 $\pm$ 5 67
TAS 340 $\pm$ 5 360
FastClick 11 $\pm$ 1 9.2
Table 5: JIT compilation time and corresponding size
Figure 10: Compilation time of LibLPM-NI (nested-if specialization) as a function of the number of elements in the LPM lookup table.
To analyze how the compilation time changes with the increase in the size of the JIT, we consider the code generated by Iridescent for the LibLPM-NI specialization. Figure 10 shows the change in the compilation time for the LibLPM-NI specialization as a function of the number of the elements in the lpm lookup table. The compilation time increases linearly with increase in the number of lpm entries. This is because Iridescent generates one basic block in the specialized code for each lpm entry. As a result, the code size generated by Iridescent grows linearly with the total number of entries.
Developer cost.
Table 6 shows the necessary lines of code changes to integrate Iridescent specializations into the target systems. For all cases, fewer than 100 lines of code change. Implementing the custom LPM specialization only required 162 lines of code for enabling the custom user-defined LPM specialization.
System Handler Annotations (LoC) Spec Policy (LoC)
MMulBlockBench 2 47
LibLPM 4 9
Network Functions 5 43
TAS 8 38
FastClick 5 88
Table 6: Lines of Code change required to integrate Iridescent.
Figure 11: Impact of sampling rate on instrumentation overhead for LibLPM-FP configuration for Workload B.
Instrumentation overhead.
To measure the overhead of adding instrumentation, we devise a microbenchmark called SimpleBench, consisting of two very simple functions, f and g. f computes the square of the input and g computes the product of its two inputs. For f, we add a general specialization point for its input a. For g, we add a range-based specialization for its second input b. For both the specialization points a and b, we first execute their respective functions, f and g, in normal mode and then use Iridescent to add instrumentation. We execute each configuration one million times. As a baseline, each function on average takes up to 8 cycles per execution in normal mode. For a, which is a general specialization point, adding instrumentation adds an overhead of around 450 to 500 cycles per execution. For b, which is a specialization point designed for more efficient instrumentation for specialization points with values in fixed ranges, adding instrumentation adds an overhead of around 1 extra cycle per operation. When possible, Iridescent users should opt to use the more efficient specialization point. However, as this ma not always be possible and users may be forced to use the general specialization point, Iridescent allows users to specify a sampling rate to avoid incurring the instrumentation cost for low-latency computations. Figure 11 shows how decreasing the sampling rate for the LibLPM-FP Workload B configuration can decrease the instrumentation overhead.
Specialization guards and failures.
To measure the cost of specialization guards and failures, we use the SimpleBench microbenchmark. For function f, we use Iridescent to specialize the value of a and generate two different versions. In the first version, we disable the specialization check at the entrypoint of f. We then execute this version of the function with the specialized value as input one million times. Each execution of this version takes 7 cycles. In the second version, we enable the specialization check at the entrypoint of f. If the specialization check fails, i.e. the input value of a does not match the specialized value of a, then this version of f throws an exception that is caught by Iridescent-inserted trampoline code which subsequently calls the generic version of the function. We first execute this version of the function with inputs such that the specialization guards always pass. This check adds an extra cycle per execution of the function. Next, we execute this version of the function with inputs such that the specialization guards always fail. This results in exception handling behavior which adds approximately 5000 cycles of overhead per execution. Thus, the raw cost of a specialization check is 1 extra cycle whereas that of a specialization failure is approximately 5000 cycles.
7 Related Work
7.1 Offline Automatic Specializations
We classify offline automatic specialization techniques in four categories. (i) Compile pass selection techniques automatically select the optimal sequence of compiler optimization passes through auto-tuning methods for trying out different combinations [62, 58, 59, 58], (ii) Specialized compiler optimization passes that can generate specialized code for specific hardware features like vectorization [56] or prefetching [73], or for other specific optimizations [47, 52, 69, 2], (iii) Feedback-Driven Optimizations (FDO) [13, 61, 34, 23, 57, 71, 79, 67, 87, 33, 90, 38, 89, 33], also known as PGO (Profile-Guided Optimizations), are compiler optimizations consisting of multiple executions of the program to be optimized by collecting relevant metrics from a benchmarking run and then using these metrics to generate optimized versions. FDO techniques in their current form lack flexibility as they are specialized to work well for only one workload (the workload it was trained on), (iv) Specialized Code Tuning approaches either use trial runs to tune specific parameters in the code [46, 64] or apply specialized optimizations to different pieces of the code [81, 64]. Unlike Iridescent, existing offline automatic optimizations cannot adapt to changes in workload or environmental settings dynamically. Consequently, these techniques are sub-optimal for the conditions they were not trained on.
7.2 Online Automatic Optimizations
We classify existing online automatic optimizations in four categories — (i) runtime code selection, (ii) configuration optimizations, (iii) generic code specializations, (iv) domain-specific specializations.
Runtime code selection.
In this optimization technique, developers pre-generate multiple different versions of the code at compile-time and then use measured runtime statistics and offline-trained models to switch between the versions [40, 76, 78, 20]. These techniques are limited as they must anticipate and generate all possible versions a priori. Consequently, in contrast to Iridescent, these techniques can not fully leverage runtime information for specialization.
Config optimizations.
Runtime auto-tuning techniques such as OtterTune [82] often perform tuning through trial sessions to find optimal values for different configuration knobs. Some AutoTuning techniques such as OPPerTune [75] and SoftSKU [77] apply the AutoTuning paradigm of measurement-driven search at runtime without any trials. However, these techniques are critically limited to environmental parameters and knobs as they work transparently with respect to the deployed application. These techniques are currently not applicable to code optimizations or specializations. Iridescent enables the auto-tuning paradigm of measurement-driven search at runtime to find the best combination of code specializations for the system.
Code specializations.
Runtime code specialization techniques [18, 17, 32, 22, 5, 65] produce more efficient versions of the code by exploiting values and invariants that only exist at runtime. Today, these specializations manifest in three ways: (i) binary rewriting techniques [3]; (ii) Transparent Dynamic Optimization (TDO) techniques such as Dynamo [6], The Transmeta Code Morphing Software [23], DynamoRIO [9]. These optimize code at runtime without requiring any modifications by capturing and optimizing traces (sequences of instructions) that are commonly executed; (iii) optimizations in JIT compilers for interpreted languages in the form of Type Specialization [15, 57, 14] for specializing the types of data for dynamically typed languages or as Value Specialization for interpreted languages [21, 49] where the parameter values of hot functions are converted into constants. These specializations may be optionally applied depending on the computation context [37, 63, 28, 54]. In contrast to Iridescent, they are generic and typically done automatically by the runtime based on internal cost models which may not be appropriate for specific application contexts. Moreover, these specializations are limited in scope and do not take the system’s end-to-end performance into consideration and do not leverage domain- and situation-specific optimizations.
Domain-specific specializations.
Recently, value specializations have extended beyond interpreted languages for specific use cases such as value specializations for GPU kernels [29], or through incremental specializations for packet processing frameworks [68, 53, 27]. Iridescent provides a general framework for configuring, applying, exploring, and selecting such specializations at runtime.
8 Discussion
Debugging.
Iridescent makes it harder for developers to debug performance issues. This is because the generated specialized code is often different from the code that was originally written by the developer, which makes it harder for the developers to fully understand the execution sequences. This is further exacerbated by the fact that different workloads may lead to different specializations making it difficult to reproduce issues seen in production. Combining Iridescent with monitoring and distributed tracing techniques could offer developers a solution to their debugging problems.
Leveraging scale & ML for exploration.
Several ML-based techniques exist for tuning configurations [82, 1, 12, 48, 50, 51, 74, 75, 26, 38, 80, 68]. These ML techniques can be implemented as algorithms with OPPerTune [75], which we have already integrated with Iridescent. Moreover, Iridescent could further incorporate SmartChoices [10] to allow different ML techniques to apply to specific specialization points. As part of future work, Iridescent could potentially also parallelize exploration across a large number of replicas to explore larger subsets of the specialization space.
Limitations.
While Iridescent is designed to operate in dynamic and irregular operating conditions, exploration overhead overshadows performance benefits if the operating conditions do not remain stable for long enough to reap the benefits of the specialization optimizations. Ideally, developers should configure the specialization policy to opt for the general version of the handler code in such scenarios.
9 Conclusions
In this work, we presented Iridescent, a framework enabling online workload-driven runtime specialization of systems to improve performance. Iridescent provides a toolkit for developers to implement system-specific specializations that utilize runtime data and invariants to generate more performant systems. Iridescent provides the necessary tooling for developers to configure automatic runtime exploration of the space of potential specializations to find the best performing configuration at each instant. By combining developer insight with automated compiler optimizations, Iridescent enables systems to adapt efficiently to changing workloads and environments with minimal manual effort.
Acknowledgments
We thank Yuchen Qian for his contributions to an early prototype Iridescent implementation.